# On the geometry of Kähler–Frobenius manifolds &\&& their classification

### v1

@@ Abstract.
The purpose of this article is to show that flat compact Kähler manifolds exhibit the structure of a Frobenius manifold, a structure originating in 2D Topological Quantum Field Theory and closely related to Joyce structure. As a result, we classify all such manifolds. It can be deduced that Kähler–Frobenius manifolds include certain Calabi–Yau manifolds, complex tori $T=\mathbb{C}^{n}/\mathbb{Z}^{n}$, Hantzsche–Wendt manifolds, hyperelliptic manifolds and manifolds of type $T/G$, where $G$ is a finite group acting on $T$ freely and containing no translations. An explicit study is provided for the two-dimensional case. Additionally, we can prove that Chern’s conjecture for Kähler pre-Frobenius manifolds holds. Lastly, we establish that certain classes of Kähler-Frobenius manifolds share a direct relationship with theta functions which are important objects in number theory as well as complex analysis.
The author acknowledges support from the project No. 2022/47/P/ST1/01177 co-founded by the National Service Centre and the European Union’s Horizon 2020 research and innovation program, under the Marie Sklodowska Curie grant agreement No. 945339.[Uncaptioned image]
@@ 1.3. Results
    Hantzsche–Wendt manifolds;
@@ 2. Frobenius bundles
  • The rank 4 tensor $(\nabla_{W}A)(X,Y,Z)$ is fully symmetric;
@@ 2.4. Frobenius manifolds: prerequisites, recollections, notations
Man is the category of analytic manifolds over a field $\mathbb{K}$ (being $\mathbb{R}$ or $\mathbb{C}$). Let $M\in Ob({\bf Man})$ be an object of Man. Denote by $\mathcal{O}_{M}$ (resp. $\Omega^{1}_{M}$) the sheaf of holomorphic (resp. holomorphic 1-forms) on $M$.
@@ 2.4.2. Affine connections
Assume $E$ is the space of sections of the tangent bundle over $M$ (in a more general context it can be a quasi coherent sheaf of $\mathcal{O}_{M}$-modules).
A connection on $E$ is a homomorphism $\nabla$ of abelian sheaves $\nabla:E\to\Omega^{1}\otimes E$ such that the Leibnitz rule is satisfied. That is
@@
where $f$ is a smooth function and $e$ is a section of $E$ defined over an open subset in $M$.
@@ 2.4.7. Metrics
Let $(E,M,\pi)$ be a fiber bundle over $M$ with structure group $G$. Assume that $M$ and the fiber space are equipped with a (non degenerate) bilinear symmetric form. For each $x\in M$, there exists a Hermitian (resp. Riemannian) inner product on the fiber $E_{x}$ defined as:
$\langle-,-\rangle:\,E_{x}\times E_{x}\to\mathbb{C}.$
For any open in $M$ and sections $\xi,\eta$ in $E$ the mapping $x\mapsto\langle\xi,\eta\rangle_{x}$ is differentiable.
@@ 2.5.1. Algebras
Assume $(\mathcal{A},\circ)$ is an algebra of rank $n$. Then, the multiplication operation:
$\circ:\mathcal{A}\times\mathcal{A}\to\mathcal{A}\quad\text{ is defined by}\quad e_{i}\circ e_{j}=\sum C_{ij}^{k}e_{k},$
where $C_{ij}^{k}$ are the structure constants of $\mathcal{A}$ and $e_{i}$ are generators of $A$.
  • The algebra $\mathcal{A}$ is commutative if the commutator vanishes:
@@
  • The algebra $\mathcal{A}$ is associative if the associator vanishes:
@@ 2.5.2. Frobenius algebra
A Frobenius algebra $(\mathcal{A},\circ)$ over $\mathbb{K}$ is an associative, commutative, unital algebra with multiplication operation ``"``"`$\circ$"` ` ∘ " equipped with a symmetric bilinear form $\langle$-,$\rangle$ - , - ⟩ satisfying
(2.5.2.1) $\langle x\circ y,z\rangle=\langle x,y\circ z\rangle,\quad\forall x,y,z\in\mathcal{A}.$
@@ 2.5.3. Pre-Frobenius algebra
A pre-Frobenius algebra is an algebra $(\mathcal{A},\circ)$ satisfying commutativity (i.e. Equation 2.5.1.1) and Equation 2.5.2.1. It is not a Frobenius algebra because it is not associative. So, we have the table in Fig.2.
@@ 2.5.4. Algebra bundle
A bundle algebra with base $M$ is a vector bundle $\mathscr{E}=(E(\mathcal{A}),M,\pi)$ endowed with a bilinear morphism $f:E\times E\to E$ such that each fiber $E_{x}=\pi^{-1}(x)$ for any $x\in M$ is endowed with the structure of the algebra $\mathcal{A}$. Given $x\in M$, if the algebra $\mathcal{A}_{x}$ has an identity element $\epsilon_{x}$ then the mapping $x\mapsto\epsilon_{x}$ is a unit section of $E$.
@@
It is assumed that in the adjoint principal bundle $P(M,\tilde{G})$ of admissible frames $R_{x}$ the matrices $\hat{S}_{b}$ (and thus structure constants of $\mathcal{A}$) are constant. For instance they have a normal Jordan form.
@@ 2.5.5. Frobenius algebra bundle
Let $M$ be a Riemannian manifold, smooth and finite dimensional. Assume $\mathscr{E}=(TM,M,\pi)$ is the tangent vector bundle. Denote by $Aff(n)$ the group of affine transformations.
@@
Let $M$ be an $n$-dimensional manifold. A Frobenius (algebra) bundle $\mathscr{E}$ is an algebra bundle with the following conditions:
  1. $\mathscr{E}$ is a flat bundle with structure group $Aff(n)$, equipped with an affine flat torsionless connection $\nabla_{0}$.
  2. $\mathscr{E}$ is a Frobenius algebra bundle, where for every $x\in M$ each fiber $E_{x}$ is a Frobenius algebra $(A_{x},\circ)$.
  3. For every $x\in M$ there exists on the fiber $E_{x}$ a Riemannian inner product
    $g=\langle-,-\rangle:\,E_{x}\times E_{x}\to\mathbb{R}.$
  4. There exists everywhere locally on $M$ a real valued potential function $\Phi$. There exists rank three symmetric tensor given in local coordinates by $C_{ijk}=\partial_{i}\partial_{j}\partial_{k}\Phi$.
  5. Structure constants of the algebra $\mathcal{A}_{x}$ are given by relations:
@@
Assume $M$ is an $n$-dimensional manifold. The manifold $M$ comes equipped with a pre-Frobenius bundle $\mathscr{E}$ if the algebra bundle is a pre-Frobenius algebra bundle.
The important difference between a pre-Frobenius bundle and a Frobenius bundle relies on the structure of the algebra $\mathcal{A}_{x}$: how structure constants are defined and on the associator ${\bf A}$. To summarise:
$\begin{cases}\forall x\in M,\quad\text{if}\quad{\bf A}=0&\text{then}\quad E_{x}\quad\text{is a Frobenius algebra};\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\par\forall x\in M,\quad\text{if}\quad{\bf A}\neq 0&\text{then}\quad E_{x}\quad\text{is a pre-Frobenius algebra}.\end{cases}$
@@ 3.1. Real case
    where $\widehat{A}$ is a global section lying in $\widehat{A}\in\Omega^{1}_{M}\otimes_{\mathcal{O}_{M}}End(T_{M})$. For an element in the endomorphism $\tilde{F}\in End(T_{M})$ and $df\in\Omega_{M}^{1}$ ($f$ is a smooth function) we get $i_{X}(df\otimes G):=Xf\cdot\tilde{F}.$ More specifically, in local flat coordinates the multiplication operation is defined as
@@ A Hermitian Frobenius manifold $M$ is a complex manifold with:
(3.2.5.4) $R_{a\bar{b}c\bar{d}}=\frac{\partial^{2}g_{a\bar{b}}}{\partial z^{c}\partial\bar{z}^{d}}-g^{e\bar{\gamma}}\frac{\partial g_{a\bar{\gamma}}}{\partial z^{c}}\frac{g_{e\bar{b}}}{\partial\bar{z}^{d}}.$
@@
$\frac{\partial^{2}g_{a\bar{b}}}{\partial z^{c}\partial\bar{z}^{d}}=g^{e\bar{\gamma}}\frac{\partial g_{a\bar{\gamma}}}{\partial z^{c}}\frac{g_{e\bar{b}}}{\partial\bar{z}^{d}}.$
@@ 4. Kähler–Frobenius manifolds
1) The manifold $M$ admits a holomorphic affine connection if and only if the 1-cocycle defining an element of $H^{1}(M,\Omega^{1}\otimes\Omega^{1}\otimes\mathcal{O})$ is zero.
This 1-cocycle is cohomologous to zero if and only if there exists a 0-coboundary in $H^{0}(U,\Omega^{1}\otimes\Omega^{1}\otimes\mathcal{O})$ such that $C_{VU}=b_{V}-b_{U}$, where
@@
(4.1.0.1) $R_{a\bar{b}c\bar{d}}=\frac{\partial^{2}g_{a\bar{b}}}{\partial z^{c}\partial\bar{z}^{d}}-g^{e\bar{\gamma}}\frac{\partial g_{a\bar{\gamma}}}{\partial z^{c}}\frac{g_{e\bar{b}}}{\partial\bar{z}^{d}}.$
@@
By construction if $M$ is a $\mathbb{C}$-Frobenius manifold then it is smooth compact flat Riemannian manifold equipped with a Kähler structure. By Johnsson [J, Theorem 4.2]: Let $M$ be a smooth compact flat Riemannian manifold. If $M$ is a flat Kähler manifold then $M$ admits also the structure of a flat smooth complex projective variety. Therefore, the conclusion follows. ∎
@@ 5.1. Kähler–Einstein compact manifolds
2)   Define the rank three (nearly-)symmetric tensor $\Phi_{ijk}=\partial_{i}\partial_{j}\partial_{\bar{k}}\Phi$.
@@
3)   We define an $\mathcal{O}_{M}$-symmetric bilinear multiplication operation $\circ$ on the tangent sheaf $T_{M}$.
@@
In the complex case, we decompose the tangent space as $\mathcal{T}_{M}=\mathcal{T}_{M}^{1,0}\oplus\mathcal{T}_{M}^{0,1}$.
So, we have a linear combination of two subalgberas defined on $\mathcal{T}_{M}^{1,0}$ and $\mathcal{T}_{M}^{0,1}$.
@@ 7. Classification of Kähler–Frobenius manifolds $n\geq 2$
A classification theorem of compact Kähler Frobenius manifolds is provided. Below the table we recall some definitions of Calabi–Yau manifolds and Hantzsche–Wendt manifolds.
@@
    Hantzsche–Wendt manifolds manifolds equipped with a rotation group,
@@
Recall a few definitions for Calabi–Yau manifolds and Hantzsche–Wendt manifolds.
  • An $n$-dimensional simply-connected compact Calabi–Yau manifold $M$ is a compact Kähler manifold with vanishing first Chern class and with $h^{0,p}=0$ for each $0<p<n$ and $n>2$. It has a trivial canonical line bundle and admits a Ricci-flat Kähler metric with holonomy group $SU(n)$.
  • A Hantzsche–Wendt $n$-manifold is a (non-necessarily) orientable flat manifold in dimension $n\geq 3$, with holonomy group $\mathbb{Z}_{2}^{n-1}$. This generalises the original version of a Hantzsche–Wendt manifold, which is a 3-dimensional manifold with vanishing first Betti number $b_{1}=0$.
The non-orientable case occurs for each dimension $n\geq 2$. The orientable case occurs only in each odd dimension $n\geq 3.$ Any complex orientable Hantzsche–Wendt manifold has a spin structure. The holonomy representation $\mathbb{Z}^{n-1}_{2}\to U(n)$ of an orientable complex Hantzsche–Wendt manifold has its image in $SU(n)$.
@@
By Theorem 7.1.0.1 one has a classification of different sources of Kähler–Frobenius manifolds. Using the computations on the number classes of flat Kähler manifolds of dimensions 2 and 3 in [DHS] we deduce that there exist 8 compact 2-dimensional Kähler–Frobenius manifolds and 174 compact 3-dimensional Kähler–Frobenius manifolds. ∎

### v2

@@ Abstract.
The purpose of this article is to show that flat compact Kähler manifolds exhibit the structure of a Frobenius manifold, a structure originating in 2D Topological Quantum Field Theory and closely related to Joyce structure. As a result, we classify all such manifolds. It can be deduced that Kähler–Frobenius manifolds include certain Calabi–Yau manifolds, complex tori $T=\mathbb{C}^{n}/\mathbb{Z}^{n}$, generalized Hantzsche–Wendt manifolds, hyperelliptic manifolds and manifolds of type $T/G$, where $G$ is a finite group acting on $T$ freely and containing no translations. An explicit study is provided for the two-dimensional case. Additionally, we can prove that Chern’s conjecture for Kähler pre-Frobenius manifolds holds. Lastly, we establish that certain classes of Kähler-Frobenius manifolds share a direct relationship with theta functions which are important objects in number theory as well as complex analysis.
The author acknowledges support from the project No. 2022/47/P/ST1/01177 co-founded by the National Science Centre and the European Union’s Horizon 2020 research and innovation program, under the Marie Sklodowska Curie grant agreement No. 945339.[Uncaptioned image]. The author thanks A. Szczepański and M.Halenda for interesting discussions about flat Kähler manifolds and for an invitation to the university of Gdańsk.
@@ 1.3. Results
    Generalized Hantzsche–Wendt manifolds (orientable);
@@ 2. Frobenius bundles
  • The rank 4 tensor $({\stackrel{{\scriptstyle\scriptscriptstyle{0}}}{{\nabla}}}\phantom{}_{W}A)(X,Y,Z)$ is fully symmetric;
@@ 2.4. Frobenius manifolds: prerequisites, recollections, notations
Man is the category of analytic manifolds over a field $\mathbb{K}$ (being $\mathbb{R}$ or $\mathbb{C}$). Let $M\in Ob({\bf Man})$ be an object of Man. Denote by $\mathscr{O}_{M}$ (resp. $\Omega^{1}_{M}$) the sheaf of holomorphic (resp. holomorphic 1-forms) on $M$.
@@ 2.4.2. Affine connections
Assume $\mathscr{E}$ is the space of sections of the tangent bundle over $M$ (in a more general context it can be a quasi coherent sheaf of $\mathscr{O}_{M}$-modules).
A connection on $\mathscr{E}$ is a homomorphism $\nabla$ of abelian sheaves $\nabla:\mathscr{E}\to\Omega^{1}\otimes\mathscr{E}$ such that the Leibnitz rule is satisfied. That is
@@
where $f$ is a smooth function and $e$ is a section of $\mathscr{E}$ defined over an open subset in $M$.
@@ 2.4.7. Metrics
Let $(\mathscr{E},M,\pi)$ be a fiber bundle over $M$ with structure group $G$. Assume that $M$ and the fiber space are equipped with a (non degenerate) bilinear symmetric form. For each $x\in M$, there exists a Hermitian (resp. Riemannian) inner product on the fiber $\mathscr{E}_{x}$ defined as:
$\langle-,-\rangle:\,\mathscr{E}_{x}\times\mathscr{E}_{x}\to\mathbb{C}.$
For any open in $M$ and sections $\xi,\eta$ in $\mathscr{E}$ the mapping $x\mapsto\langle\xi,\eta\rangle_{x}$ is differentiable.
@@ 2.5.1. Algebras
Assume $(\mathscr{A},\circ)$ is an algebra of rank $n$. Then, the multiplication operation:
$\circ:\mathscr{A}\times\mathscr{A}\to\mathscr{A}\quad\text{is defined by}\quad e_{i}\circ e_{j}=\sum_{k}C_{ij}^{k}e_{k},$
where $C_{ij}^{k}$ are the structure constants of $\mathscr{A}$ and $e_{i}$ are generators of $A$.
  • The algebra $\mathscr{A}$ is commutative if the commutator vanishes:
@@
  • The algebra $\mathscr{A}$ is associative if the associator vanishes:
@@ 2.5.2. Frobenius algebra
A Frobenius algebra $(\mathscr{A},\circ)$ over $\mathbb{K}$ is an associative, commutative, unital algebra with multiplication operation ``"``"`$\circ$"` ` ∘ " equipped with a symmetric bilinear form $\langle$-,$\rangle$ - , - ⟩ satisfying
(2.5.2.1) $\langle x\circ y,z\rangle=\langle x,y\circ z\rangle,\quad\forall x,y,z\in\mathscr{A}.$
@@ 2.5.3. Pre-Frobenius algebra
A pre-Frobenius algebra is an algebra $(\mathscr{A},\circ)$ satisfying commutativity (i.e. Equation 2.5.1.1) and Equation 2.5.2.1. It is not a Frobenius algebra because it is not associative. So, we have the table in Fig.2.
@@ 2.5.4. Algebra bundle
A bundle algebra with base $M$ is a vector bundle ${\pazocal{E}}=(\mathscr{E}(\mathscr{A}),M,\pi)$ endowed with a bilinear morphism such that each fiber $\mathscr{E}_{x}=\pi^{-1}(x)$ for any $x\in M$ is endowed with the structure of the algebra $\mathscr{A}$. Given $x\in M$, if the algebra $\mathscr{A}_{x}$ has an identity element $\epsilon_{x}$ then the mapping $x\mapsto\epsilon_{x}$ is a unit section of $\mathscr{E}$.
@@
It is assumed that in the adjoint principal bundle $P(M,\tilde{G})$ of admissible frames $R_{x}$ the matrices $\hat{S}_{b}$ (and thus structure constants of $\mathscr{A}$) are constant. For instance they have a normal Jordan form.
@@ 2.5.5. Frobenius algebra bundle
Let $M$ be a Riemannian manifold, smooth and finite dimensional. Assume $\pazocal{E}=(TM,M,\pi)$ is the tangent vector bundle. Denote by $Aff(n)$ the group of affine transformations.
@@
Let $M$ be an $n$-dimensional manifold. A Frobenius (algebra) bundle $\pazocal{E}$ is an algebra bundle with the following conditions:
  1. $\pazocal{E}$ is a flat bundle with structure group $Aff(n)$, equipped with an affine flat torsionless connection ${\stackrel{{\scriptstyle\scriptscriptstyle{0}}}{{\nabla}}}\phantom{}$.
  2. $\pazocal{E}$ is a Frobenius algebra bundle, where for every $x\in M$ each fiber $\mathscr{E}_{x}$ is a Frobenius algebra $(A_{x},\circ)$.
  3. For every $x\in M$ there exists on the fiber $\mathscr{E}_{x}$ a Riemannian inner product
    $g=\langle-,-\rangle:\,\mathscr{E}_{x}\times\mathscr{E}_{x}\to\mathbb{R}.$
  4. There exists everywhere locally on $M$ a real valued potential function $\Phi$. There exists rank three symmetric tensor given in local coordinates by $C_{ijk}=\partial_{i}\partial_{j}\partial_{k}\Phi$ where $\partial_{l}=\frac{\partial}{\partial x^{l}}$.
  5. Structure constants of the algebra $\mathscr{A}_{x}$ are given by relations:
@@
Assume $M$ is an $n$-dimensional manifold. The manifold $M$ comes equipped with a pre-Frobenius bundle $\pazocal{E}$ if the algebra bundle is a pre-Frobenius algebra bundle.
The important difference between a pre-Frobenius bundle and a Frobenius bundle relies on the structure of the algebra $\mathscr{A}_{x}$: how structure constants are defined and on the associator ${\bf A}$. To summarise:
$\begin{cases}\forall x\in M,\quad\text{if}\quad{\bf A}=0&\text{then}\quad\mathscr{E}_{x}\quad\text{is a Frobenius algebra};\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\par\forall x\in M,\quad\text{if}\quad{\bf A}\neq 0&\text{then}\quad\mathscr{E}_{x}\quad\text{is a pre-Frobenius algebra}.\end{cases}$
@@ 3.1. Real case
    where $\widehat{A}$ is a global section lying in $\widehat{A}\in\Omega^{1}_{M}\otimes_{\mathscr{O}_{M}}End(T_{M})$. For an element in the endomorphism $\tilde{F}\in End(T_{M})$ and $df\in\Omega_{M}^{1}$ ($f$ is a smooth function) we get $i_{X}(df\otimes G):=Xf\cdot\tilde{F}.$ More specifically, in local flat coordinates the multiplication operation is defined as
@@ A Hermitian Frobenius manifold $M$ is a complex manifold with:
(3.2.5.4) $R_{a\bar{b}c\bar{d}}=\frac{\partial^{2}g_{a\bar{b}}}{\partial z^{c}\partial\bar{z}^{d}}-g^{e\bar{\gamma}}\frac{\partial g_{a\bar{\gamma}}}{\partial z^{c}}\frac{\partial g_{e\bar{b}}}{\partial\bar{z}^{d}}.$
@@
$\frac{\partial^{2}g_{a\bar{b}}}{\partial z^{c}\partial\bar{z}^{d}}=g^{e\bar{\gamma}}\frac{\partial g_{a\bar{\gamma}}}{\partial z^{c}}\frac{\partial g_{e\bar{b}}}{\partial\bar{z}^{d}}.$
@@ 4. Kähler–Frobenius manifolds
1) The manifold $M$ admits a holomorphic affine connection if and only if the 1-cocycle defining an element of $H^{1}(M,\Omega^{1}\otimes\Omega^{1}\otimes\mathscr{O})$ is zero.
This 1-cocycle is cohomologous to zero if and only if there exists a 0-coboundary in $H^{0}(U,\Omega^{1}\otimes\Omega^{1}\otimes\mathscr{O})$ such that $C_{VU}=b_{V}-b_{U}$, where
@@
(4.1.0.1) $R_{a\bar{b}c\bar{d}}=\frac{\partial^{2}g_{a\bar{b}}}{\partial z^{c}\partial\bar{z}^{d}}-g^{e\bar{\gamma}}\frac{\partial g_{a\bar{\gamma}}}{\partial z^{c}}\frac{\partial g_{e\bar{b}}}{\partial\bar{z}^{d}}.$
@@
By construction if $M$ is a $\mathbb{C}$-Frobenius manifold then it is smooth compact flat Riemannian manifold equipped with a Kähler structure. By Johnsson [J, Theorem 4.2], we have the following. Let $M$ be a smooth compact flat Riemannian manifold. If $M$ is a flat Kähler manifold then $M$ admits also the structure of a flat smooth complex projective variety. Therefore, the conclusion follows. ∎
@@ 5.1. Kähler–Einstein compact manifolds
2)   Define the rank three symmetric tensor $\Phi_{ijk}=\partial_{i}\partial_{j}\partial_{\bar{k}}\Phi$.
@@
3)   We define an $\mathscr{O}_{M}$-symmetric bilinear multiplication operation $\circ$ on the tangent sheaf $T_{M}$.
@@
In the complex case, we decompose the tangent space as $\mathscr{T}_{M}=\mathscr{T}_{M}^{1,0}\oplus\mathscr{T}_{M}^{0,1}$.
So, we have a linear combination of two subalgberas defined on $\mathscr{T}_{M}^{1,0}$ and $\mathscr{T}_{M}^{0,1}$.
@@ 7. Classification of Kähler–Frobenius manifolds $n\geq 2$
A classification theorem of compact Kähler Frobenius manifolds is provided. Below the table we recall some definitions of Calabi–Yau manifolds and generalized Hantzsche–Wendt manifolds.
@@
    General. Hantzsche–Wendt manifolds manifolds equipped with a rotation group,
@@
Recall the definition of a generalized Hantzsche–Wendt manifolds.
A Hantzsche–Wendt $n$-manifold is a (non-necessarily) orientable flat manifold in dimension $n\geq 3$, with holonomy group $\mathbb{Z}_{2}^{n-1}$. This generalises the original version of a Hantzsche–Wendt manifold, which is a 3-dimensional manifold with vanishing first Betti number $b_{1}=0$. The non-orientable case occurs for each dimension $n\geq 2$. The orientable case occurs only in each odd dimension $n\geq 3.$ Any complex orientable Hantzsche–Wendt manifold has a spin structure. The holonomy representation $\mathbb{Z}^{n-1}_{2}\to U(n)$ of an orientable complex Hantzsche–Wendt manifold has its image in $SU(n)$.
@@
By Theorem 7.1.0.1 one has a classification of different sources of Kähler–Frobenius manifolds. Using the computations on the number classes of flat Kähler manifolds of dimensions 2 and 3 and the classification criterion in [DHS] we deduce that there exist 8 compact 2-dimensional Kähler–Frobenius manifolds and 174 compact 3-dimensional Kähler–Frobenius manifolds. ∎